OpenStudy (anonymous):

Help With One Question!! Please do "NOT" tell me the answer Thank Youu! You deposit $2,500 in an account that pays 5.7% interest compounded annually. What is the balance after five years?

OpenStudy (hoblos):

do you know the formula?

OpenStudy (amistre64):

at the end of each year you take the interest collected on the interest and add it back into the account. \[a_n=I\ a_{n-1}+a_{n-1}\]

OpenStudy (amistre64):

the interest collected on the account ...

OpenStudy (anonymous):

to increase a number by 5.7% is to increase it by 0.057 if you call the number \(P\) then to increase it by 0.057 means you will have \[P+.057P\] which, by the distributive property is the same as \(P(1.057)\)

OpenStudy (anonymous):

if you increase it by .057 twice, that is the same as \(P(1.057)\times (1.057)=P(1.057)^2\)

OpenStudy (anonymous):

now you want to increase it by .057 5 times, so you should be good to go right?

OpenStudy (anonymous):

i multiplied .057 5 times and i got a huge number

OpenStudy (amistre64):

what is $2,500*.057 equal ?

OpenStudy (amistre64):

yr1: 2500 * .057 added back to 2500 balance = (2500*.057 + 2500) yr2: (2500*.057 + 2500)*.057 added back to (2500*.057 + 2500) balance = (2500*.057 + 2500)*.057 + (2500*.057 + 2500) this is the basic workings of it, and we can simplify the process if we want to

OpenStudy (anonymous):

Thanks I got It now :)

OpenStudy (amistre64):

yep, good luck :)